OpenStudy (anonymous):

Write the complex number in standard form: 6 + sqrt of -4

OpenStudy (anonymous):

6+2i

OpenStudy (anonymous):

So all I would have to do with square roots is simplify them? How did you get i? I know i = 1

OpenStudy (anonymous):

or is it because of a + bi?

OpenStudy (anonymous):

pretty much...i is an imaginary number, so its the negative square root. cuz the square root of -4 doesnt exist, therefore it must be i

zepdrix (zepdrix):

\[\large i=\sqrt{-1}\] \[\large 6+\sqrt{-4}\qquad = \qquad 6+\sqrt{4}\cdot \sqrt{-1}\qquad =\qquad 6+\sqrt4 i\]

OpenStudy (anonymous):

so I keep the square root over 4?

zepdrix (zepdrix):

no c: You should be able simplify that a bit further, i was just trying to explain where the i came from.

OpenStudy (anonymous):

oh okay! and then you'd just simplify down to 2i..I understand, thanks!